OpenStudy (anonymous):

medal and fan award What is the value of this expression? e^(-4/3)

OpenStudy (anonymous):

when you have negative exponent, the folowing rule is aplyed: a^-b=1/a^b when exponent is a fraction: a^1/2=sqrt(a), a^1/3= cubicroot (a) and so on